OpenStudy (anonymous):

Help please !!!!!!!:) **** A container ship traveled to Guam and back. It took one hour longer to go there than it did to come back. The average speed on the trip was 20km/h. The average speed on the way back was 25km/h. How many hours did the trip there take? Show work or explain. Thank You!!! XD <3

OpenStudy (raden):

the distance to go = distance to come back d1 = d2 we knowed that the formula for distance (d) = speed(s) * time(t) therefore, take an equation : d1 = d2 s1 (t+1) = s2 t (given s1 = 20 km/h and s2 = 25 km/h) thus s1 (t+1) = s2 t 20(t + 1) = 25t solve fot t = ?

OpenStudy (raden):

got it ?